OpenStudy (anonymous):

The perimeter of a square is 68 meters. The side length can be represented as (8x - 3) meters. What is the value of x? (1 points) 2.5 8.875 8.125 1.75

OpenStudy (anonymous):

\(P = 4a\) Where 'a' is the side. Plug in what we know: \(68 = 4(8x - 3)\) Distribute 4 into the parenthesis: \(68 = 32x - 12\) Add 12 to both sides, what's 68 + 12?

OpenStudy (anonymous):

Do this one the same as the last problem. Use the side length for both L and W. You can also use the equation P = 4S instead

OpenStudy (anonymous):

80

OpenStudy (anonymous):

68 = 4 (8x = 3) 68 = 32x - 12 68 + 12 = 32x 80 = 32x

OpenStudy (anonymous):

Yes, that gives us: 80 = 32x Now we divide 32 to both sides, what's 80 / 32?

OpenStudy (anonymous):

x = 80/32

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

x = 2.5

OpenStudy (anonymous):

CORRECT!

OpenStudy (anonymous):

Yep, you got it.